# Stochastic theta methods for random periodic solution of stochastic differential equations under non-globally Lipschitz conditions11footnotemark: 1

### v1

@@ 1 Introduction
$\text{d}X_{t}^{t_{0}}=\big{(}-AX_{t}^{t_{0}}+f(t,X_{t}^{t_{0}})\big{)}{\,\rm{d}}{t}+g(t,X^{t_{0}}_{t}){\,\rm{d}}{W_{t}},\quad t\geq t_{0}$ (1.1)
with $X_{t_{0}}^{t_{0}}=\xi$, where $f\colon\mathbb{R}\times\mathbb{R}^{d}\to\mathbb{R}^{d}$, $g\colon\mathbb{R}\times\mathbb{R}^{d}\to\mathbb{R}^{d\times m}$ are continuous functions and $A$ is a positive definite $d\times d$ symmetric matrix. Besides, $\{W_{t}\}_{t\in\mathbb{R}}$ is a standard two-sided Wiener process in $\mathbb{R}^{m}$ on a probability space $(\Omega,\mathcal{F},\mathbb{P})$. Here the notation $X_{t}^{t_{0}},t\geq t_{0}$ denotes the solution process starting from $t_{0}$. Under Assumption 2.1 specified later, (1.1) admits a unique solution $\{X_{t}^{t_{0}}\}_{t\geq t_{0}}$ (we also rewrite it as $\{X_{t}^{t_{0}}(\xi)\}_{t\geq t_{0}}$ to emphasize its dependence on the initial value $\xi$), described by the variation of constant formula as follows
$X_{t}^{t_{0}}(\xi)=e^{-A(t-t_{0})}\xi+\int_{t_{0}}^{t}e^{-A(t-s)}f(s,X_{s}^{t_{0}}){\,\rm{d}}{s}+\int_{t_{0}}^{t}e^{-A(t-s)}g(s,X_{s}^{t_{0}}){\,\rm{d}}{W_{s}},\quad t\geq t_{0},$ (1.2)
Then the map $u(t,t_{0})\colon\Omega\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ defined by $u(t,t_{0})\xi=X_{t}^{t_{0}}(\xi)$ satisfies the semi-flow property $u(t,r,\omega)=u(t,s,\omega)\circ u(s,r,\omega)$ and the periodic property (with period $\tau>0$) $u(t+\tau,s+\tau,\omega)=u(t,s,\Theta_{t}\omega)$ with $\Theta_{t}(\omega)(s):=W_{t+s}-W_{t}$ for all $s\leq r\leq t,\omega\in\Omega$. Theorem 2.4 shows that the pull-back sequence $\{X_{t}^{-k\tau}(\xi)\}_{k\in\mathbb{N}}$ has a limit $X_{t}^{*}$ in $L^{2}(\Omega;\mathbb{R}^{d})$, where $X_{t}^{*}$ is the random periodic solution of (1.1) and is given by
@@
for all $j\in N$, where $\hat{X}_{-k\tau}^{-k\tau}=\xi$ is the initial value and $\Delta W_{-k\tau+t_{j}}:=W_{-k\tau+t_{j+1}}-W_{-k\tau+t_{j}}$ is the Brownian increment. Based on the uniform bounds for the second moment of the numerical approximation $\{\hat{X}_{-k\tau+t_{j}}^{-k\tau}\}_{j\in N}$, we employ Lemma 3.2 to show that each ST method admits a unique random periodic solution $\hat{X}_{t}^{*}$. Without resorting to the discrete Gronwall inequality, a technical iterative argument is introduced in the error estimate to overcome the difficulties arising by the superlinear growth of coefficients and the unboundedness of the time intervals. With the help of the Hölder continuity of the exact solution, we derive a uniform bound for the error between $\hat{X}_{-k\tau+t_{j}}^{-k\tau}$ and $X_{-k\tau+t_{j}}^{-k\tau}$ for all $k,j\in\mathbb{N}$, and consequently establish the mean square convergence of random periodic solutions. To be precise, the mean square convergence order of each ST method is $1/2$ for SDEs with multiplicative noise and $1$ for SDEs with additive noise, which, in particular, coincides with [7, Theorem 4.6] investigating ST method with $\theta=1$ for SDEs with additive noise.
@@ 2 Assumptions and random periodic solutions of SDEs
Suppose the following conditions are satisfied. (A1) The eigenvalues $\{\lambda_{i}\}_{i=1,2,\cdots,d}$ of the symmetric matrix $A$ satisfy $\lambda:=\min\{\lambda_{i}\}_{i=1,2,\cdots,d}>0.$ (A2) The coefficients $f$ and $g$ are continuous and periodic in time with period $\tau>0$, i.e.,
@@
(A3) There exist constants $\gamma\geq 1$, $p^{*}>4\gamma-2$, $L_{f}\in(0,\lambda)$ and $C>0$ such that for all $x,y\in\mathbb{R}^{d}$ and $t,s\in\mathbb{R}$,
$\begin{split}&~{}\langle x-y,f(t,x)-f(s,y)\rangle+(p^{*}-1)|g(t,x)-g(s,y)|^{2}\\ \leq&~{}L_{f}|x-y|^{2}+C(1+|x|+|y|)^{\gamma+1}|t-s|,\quad x,y\in\mathbb{R}^{d},t,s\in\mathbb{R},\end{split}$ (2.2)
@@
(A4) For any $p>0$, there exists a constant $C>0$ such that $|\xi|_{L^{p}(\Omega;\mathbb{R}^{d})}\leq C$.
@@
According to (2.3) and (2.7), we know that $f$ and $g$ are locally Lipschitz continuous. In combination with the monotone condition (2), [14, Theorem 3.5 in Chapter 2] ensures that (1.1) admits a unique strong solution $\{X_{t}^{t_{0}}\}_{t\geq t_{0}}$, given by
$X_{t}^{t_{0}}=\xi+\int_{t_{0}}^{t}-AX_{s}^{t_{0}}+f(s,X_{s}^{t_{0}}){\,\rm{d}}{s}+\int_{t_{0}}^{t}g(s,X_{s}^{t_{0}}){\,\rm{d}}{W_{s}},\quad t\geq t_{0}.$ (2.9)
To establish the existence and uniqueness of random periodic solution of (1.1), two auxiliary lemmas are stated and proved below. The first one considers the uniform boundedness for the exact solution of (1.1).
@@
Applying Itô formula shows that for any $k\in\mathbb{N}$, $t\geq-k\tau$ and $p\in(2,p^{*}]$,
$e^{p\lambda t}|X_{t}^{-k\tau}|^{p}=$ $~{}e^{-p\lambda k\tau}|\xi|^{p}+\int_{-k\tau}^{t}p\lambda e^{p\lambda s}|X_{s}^{-k\tau}|^{p}+pe^{p\lambda s}|X_{s}^{-k\tau}|^{p-2}\big{\langle}X_{s}^{-k\tau},-AX_{s}^{-k\tau}\big{\rangle}$
$~{}+pe^{p\lambda s}|X_{s}^{-k\tau}|^{p-2}\big{\langle}X_{s}^{-k\tau},f(s,X_{s}^{-k\tau})\big{\rangle}+\frac{p}{2}e^{p\lambda s}|X_{s}^{-k\tau}|^{p-2}|g(s,X_{s}^{-k\tau})|^{2}$
$~{}+\frac{p(p-2)}{2}e^{p\lambda s}|X_{s}^{-k\tau}|^{p-4}|(X_{s}^{-k\tau})^{\top}g(s,X_{s}^{-k\tau})|^{2}{\,\rm{d}}{s}$
$~{}+\int_{-k\tau}^{t}pe^{p\lambda s}|X_{s}^{-k\tau}|^{p-2}\big{\langle}X_{s}^{-k\tau},g(s,X_{s}^{-k\tau}){\,\rm{d}}{W_{s}}\big{\rangle}.$
@@
$~{}\mathbb{E}\big{[}e^{p\lambda t}|X_{t}^{-k\tau}|^{p}\big{]}$
$\leq$ $~{}\mathbb{E}\big{[}e^{-p\lambda k\tau}|\xi|^{p}\big{]}+\int_{-k\tau}^{t}pe^{p\lambda s}\mathbb{E}\big{[}|X_{s}^{-k\tau}|^{p-2}\big{(}\lambda|X_{s}^{-k\tau}|^{2}+\big{\langle}X_{s}^{-k\tau},-AX_{s}^{-k\tau}\big{\rangle}\big{)}\big{]}{\,\rm{d}}{s}$
$~{}+\int_{-k\tau}^{t}pe^{p\lambda s}\mathbb{E}\Big{[}|X_{s}^{-k\tau}|^{p-2}\Big{(}\big{\langle}X_{s}^{-k\tau},f(s,X_{s}^{-k\tau})\big{\rangle}+\frac{p-1}{2}|g(s,X_{s}^{-k\tau})|^{2}\Big{)}\Big{]}{\,\rm{d}}{s}$
$\leq$ $~{}\mathbb{E}\big{[}|\xi|^{p}\big{]}+\int_{-k\tau}^{t}pe^{p\lambda s}\big{(}\widetilde{L_{f}}\mathbb{E}\big{[}|X_{s}^{-k\tau}|^{p}\big{]}+\widetilde{C}\mathbb{E}\big{[}|X_{s}^{-k\tau}|^{p-2}\big{]}\big{)}{\,\rm{d}}{s}.$
@@
$\widetilde{C}|X_{s}^{-k\tau}|^{p-2}=$ $~{}\frac{p}{p-2}\times\bigg{(}(\lambda-\widetilde{L_{f}}-\varepsilon)^{\frac{p-2}{p}}|X_{s}^{-k\tau}|^{p-2}\times\frac{p-2}{p}(\lambda-\widetilde{L_{f}}-\varepsilon)^{\frac{2-p}{p}}\widetilde{C}\bigg{)}$
$\leq$ $~{}\frac{p}{p-2}\times\bigg{(}\frac{p-2}{p}(\lambda-\widetilde{L_{f}}-\varepsilon)|X_{s}^{-k\tau}|^{p}+\frac{2}{p}\Big{(}\frac{p-2}{p}(\lambda-\widetilde{L_{f}}-\varepsilon)^{\frac{2-p}{p}}\widetilde{C}\Big{)}^{\frac{p}{2}}\bigg{)}$
$\leq$ $~{}(\lambda-\widetilde{L_{f}}-\varepsilon)|X_{s}^{-k\tau}|^{p}+\frac{2}{p-2}\big{(}(\lambda-\widetilde{L_{f}}-\varepsilon)^{\frac{2-p}{p}}\widetilde{C}\big{)}^{\frac{p}{2}},$
@@
$\mathbb{E}\big{[}e^{p\lambda t}|X_{t}^{-k\tau}|^{p}\big{]}\leq\mathbb{E}\big{[}|\xi|^{p}\big{]}+Ce^{p\lambda t}+\int_{-k\tau}^{t}p(\lambda-\varepsilon)\mathbb{E}\big{[}e^{p\lambda s}|X_{s}^{-k\tau}|^{p}\big{]}{\,\rm{d}}{s}.$
@@
$\mathbb{E}\big{[}e^{p\lambda t}|X_{t}^{-k\tau}|^{p}\big{]}\leq$ $~{}\mathbb{E}\big{[}|\xi|^{p}\big{]}+Ce^{p\lambda t}+\int_{-k\tau}^{t}p(\lambda-\varepsilon)\big{(}\mathbb{E}\big{[}|\xi|^{p}\big{]}+Ce^{p\lambda s}\big{)}e^{\int_{s}^{t}p(\lambda-\varepsilon){\,\rm{d}}{u}}{\,\rm{d}}{s}$
$\leq$ $~{}\mathbb{E}\big{[}|\xi|^{p}\big{]}+Ce^{p\lambda t}+\mathbb{E}\big{[}|\xi|^{p}\big{]}e^{p(\lambda-\varepsilon)(t+k\tau)}+C\frac{\lambda-\varepsilon}{\varepsilon}e^{p\lambda t}.$
@@
Let $\{X_{t}^{-k\tau}\}_{t\geq-k\tau}$ and $\{Y_{t}^{-k\tau}\}_{t\geq-k\tau}$ be two solutions of (1.1) with different initial values $\xi$ and $\eta$, respectively. Suppose that Assumption 2.1 holds for both initial values $\xi$ and $\eta$. Then for any $p\in[2,p^{*}],k\in\mathbb{N}$ and $t\geq-k\tau$,
$\mathbb{E}\big{[}|X_{t}^{-k\tau}-Y_{t}^{-k\tau}|^{p}\big{]}\leq e^{p(L_{f}-\lambda)(t+k\tau)}\mathbb{E}\big{[}|\xi-\eta|^{p}\big{]}.$ (2.11)
@@
$~{}e^{p\lambda t}|X_{t}^{-k\tau}-Y_{t}^{-k\tau}|^{p}$
$=$ $~{}e^{-p\lambda k\tau}|\xi-\eta|^{p}+\int_{-k\tau}^{t}p\lambda e^{p\lambda s}|X_{s}^{-k\tau}-Y_{s}^{-k\tau}|^{p}$
$~{}+pe^{p\lambda s}|X_{s}^{-k\tau}-Y_{s}^{-k\tau}|^{p-2}\big{\langle}X_{s}^{-k\tau}-Y_{s}^{-k\tau},-A(X_{s}^{-k\tau}-Y_{s}^{-k\tau})\big{\rangle}{\,\rm{d}}{s}$
$~{}+\int_{-k\tau}^{t}pe^{p\lambda s}|X_{s}^{-k\tau}-Y_{s}^{-k\tau}|^{p-2}\big{\langle}X_{s}^{-k\tau}-Y_{s}^{-k\tau},f(s,X_{s}^{-k\tau})-f(s,Y_{s}^{-k\tau})\big{\rangle}$
$~{}+\frac{p}{2}e^{p\lambda s}|X_{s}^{-k\tau}-Y_{s}^{-k\tau}|^{p-2}|g(s,X_{s}^{-k\tau})-g(s,Y_{s}^{-k\tau})|^{2}$
$~{}+\frac{p(p-2)}{2}e^{p\lambda s}|X_{s}^{-k\tau}-Y_{s}^{-k\tau}|^{p-4}|(X_{s}^{-k\tau}-Y_{s}^{-k\tau})^{\top}g(s,X_{s}^{-k\tau})-g(s,Y_{s}^{-k\tau})|^{2}{\,\rm{d}}{s}$
$~{}+\int_{-k\tau}^{t}pe^{p\lambda s}|X_{s}^{-k\tau}-Y_{s}^{-k\tau}|^{p-2}\big{\langle}X_{s}^{-k\tau}-Y_{s}^{-k\tau},(g(s,X_{s}^{-k\tau})-g(s,Y_{s}^{-k\tau})){\,\rm{d}}{W_{s}}\big{\rangle}{\,\rm{d}}{s}.$
@@
$\mathbb{E}\big{[}e^{p\lambda t}|X_{t}^{-k\tau}-Y_{t}^{-k\tau}|^{p}\big{]}\leq$ $~{}\mathbb{E}\big{[}e^{-p\lambda k\tau}|\xi-\eta|^{p}\big{]}+\int_{-k\tau}^{t}p\mathbb{E}\Big{[}e^{p\lambda s}|X_{s}^{-k\tau}-Y_{s}^{-k\tau}|^{p-2}$
×$~{}\times\Big{(}\big{\langle}X_{s}^{-k\tau}-Y_{s}^{-k\tau},f(s,X_{s}^{-k\tau})-f(s,Y_{s}^{-k\tau})\big{\rangle}$
$~{}+\frac{p-1}{2}|g(s,X_{s}^{-k\tau})-g(s,Y_{s}^{-k\tau})|^{2}\Big{)}\Big{]}{\,\rm{d}}{s}$
$\leq$ $~{}\mathbb{E}\big{[}e^{-p\lambda k\tau}|\xi-\eta|^{p}\big{]}+\int_{-k\tau}^{t}pL_{f}\mathbb{E}\big{[}e^{p\lambda s}|X_{s}^{-k\tau}-Y_{s}^{-k\tau}|^{p}\big{]}{\,\rm{d}}{s}.$
@@
$\mathbb{E}\big{[}e^{p\lambda t}|X_{t}^{-k\tau}-Y_{t}^{-k\tau}|^{p}\big{]}\leq$ $~{}\mathbb{E}\big{[}e^{-p\lambda k\tau}|\xi-\eta|^{p}\big{]}e^{pL_{f}(t+k\tau)},$

### v2

@@ 1 Introduction
$\text{d}X_{t}^{t_{0}}=\big{(}-AX_{t}^{t_{0}}+f(t,X_{t}^{t_{0}})\big{)}{\,\rm{d}}{t}+g(t,X^{t_{0}}_{t}){\,\rm{d}}{W_{t}},\quad t>t_{0}$ (1.1)
where $X_{t_{0}}^{t_{0}}=\xi$ is a $\mathbb{R}^{d}$-valued random variable, $f\colon\mathbb{R}\times\mathbb{R}^{d}\to\mathbb{R}^{d}$, $g\colon\mathbb{R}\times\mathbb{R}^{d}\to\mathbb{R}^{d\times m}$ are continuous functions and $A$ is a positive definite $d\times d$ symmetric matrix. Besides, $\{W_{t}\}_{t\in\mathbb{R}}$ is a standard two-sided Wiener process in $\mathbb{R}^{m}$ on a probability space $(\Omega,\mathcal{F},\mathbb{P})$. Here the notation $X_{t}^{t_{0}},t\geq t_{0}$ denotes the solution process starting from $t_{0}$. Under Assumption 2.1 specified later, (1.1) admits a unique solution $\{X_{t}^{t_{0}}\}_{t\geq t_{0}}$ (we also rewrite it as $\{X_{t}^{t_{0}}(\xi)\}_{t\geq t_{0}}$ to emphasize its dependence on the initial value $\xi$), described by the variation of constant formula as follows
$X_{t}^{t_{0}}(\xi)=e^{-A(t-t_{0})}\xi+\int_{t_{0}}^{t}e^{-A(t-s)}f(s,X_{s}^{t_{0}}){\,\rm{d}}{s}+\int_{t_{0}}^{t}e^{-A(t-s)}g(s,X_{s}^{t_{0}}){\,\rm{d}}{W_{s}},\quad t\geq t_{0}.$ (1.2)
Then the map $u_{t_{0},t}\colon\Omega\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ defined by $u_{t_{0},t}(\omega,\xi)=X_{t}^{t_{0}}(\omega,\xi)$ satisfies the semiflow property $u_{r,t}(\omega,\xi)=u_{s,t}(\omega,u_{r,s}(\omega,\xi))$ and the periodic property (with period $\tau>0$) $u_{s+\tau,t+\tau}(\omega,\xi)=u_{s,t}(\Theta_{\tau}\omega,\xi)$ with $\Theta_{t}(\omega)(s):=W_{t+s}-W_{t}$ for all $r\leq s\leq t,\omega\in\Omega$ and $\xi\in\mathbb{R}^{d}$. Theorem 2.6 shows that the pull-back sequence $\{X_{t}^{-k\tau}(\xi)\}_{k\in\mathbb{N}}$ has a limit $X_{t}^{*}$ in $L^{2}(\Omega;\mathbb{R}^{d})$, where $X_{t}^{*}$ is the random periodic solution of (1.1) and is given by
@@
for all $j\in\mathbb{N}$, where $\hat{X}_{-k\tau}^{-k\tau}=\xi$ is the initial value and $\Delta W_{-k\tau+t_{j}}:=W_{-k\tau+t_{j+1}}-W_{-k\tau+t_{j}}$ is the Brownian increment. Based on the uniform bounds for the second moment of the numerical approximation $\{\hat{X}_{-k\tau+t_{j}}^{-k\tau}\}_{j\in\mathbb{N}}$, we employ Lemma 3.2 to show that each ST method admits a unique random periodic solution $\hat{X}_{t}^{*}$. Without resorting to the discrete Gronwall inequality, a technical iterative argument is introduced in the error estimate to overcome the difficulties arising by the superlinear growth of coefficients and the unboundedness of the time intervals. With the help of the Hölder continuity of the exact solution, we derive a uniform bound for the error between $\hat{X}_{-k\tau+t_{j}}^{-k\tau}$ and $X_{-k\tau+t_{j}}^{-k\tau}$ for all $k,j\in\mathbb{N}$, and consequently establish the mean square convergence of random periodic solutions. To be precise, the mean square convergence order of each ST method is $1/2$ for SDEs with multiplicative noise and $1$ for SDEs with additive noise, which, in particular, coincides with [7, Theorem 4.6] investigating ST method with $\theta=1$ for SDEs with additive noise.
@@ 2 Assumptions and random periodic solutions of SDEs
Suppose the following conditions are satisfied.
(A1) The eigenvalues $\{\lambda_{i}\}_{i=1,2,\cdots,d}$ of the symmetric matrix $A$ satisfy $\lambda:=\min\{\lambda_{i}\}_{i=1,2,\cdots,d}>0.$
(A2) The coefficients $f$ and $g$ are continuous and periodic in time with period $\tau>0$, i.e.,
@@
(A3) There exist constants $\gamma\geq 1$, $p^{*}>(d+4)\gamma$, $L_{f}\in(0,\lambda)$ and $C>0$ such that for all $x,y\in\mathbb{R}^{d}$ and $t,s\in\mathbb{R}$,
$\langle x-y,f(t,x)-f(s,y)\rangle+(p^{*}-1)|g(t,x)-g(s,y)|^{2}\leq L_{f}|x-y|^{2}+C(1+|x|+|y|)^{\gamma+1}|t-s|$ (2.2)
@@
(A4) For any $0<p\leq p^{*}$, there exists a constant $C>0$ such that $|\xi|_{L^{p}(\Omega;\mathbb{R}^{d})}\leq C$.
(A5) Let $\Pi(s,x)$ be the explosion time of the solution $X_{t}^{s}(x)$ to (1.1) with initial value $x\in\mathbb{R}^{d}$ and satisfy $\mathbb{P}(\Pi(s,x)=\infty,s\in\mathbb{R},x\in\mathbb{R}^{d})=1$.
@@
According to (2.3) and (2.7), we know that $f$ and $g$ are locally Lipschitz continuous on $[t_{0},T]\times\mathbb{R}^{d}$, where $[t_{0},T]$ is a finite subinterval of $[t_{0},\infty)$ for any $T>t_{0}$. In combination with the monotone condition (2), [14, Theorem 3.5 in Chapter 2] ensures that (1.1) admits a unique strong solution $\{X_{t}^{t_{0}}\}_{t\in[t_{0},T]}$, given by
$X_{t}^{t_{0}}=\xi+\int_{t_{0}}^{t}-AX_{s}^{t_{0}}+f(s,X_{s}^{t_{0}}){\,\rm{d}}{s}+\int_{t_{0}}^{t}g(s,X_{s}^{t_{0}}){\,\rm{d}}{W_{s}},\quad t\in[t_{0},T].$ (2.9)
Due to the arbitrariness of $[t_{0},T]\subset[t_{0},\infty)$, (1.1) possesses a unique global solution $\{X_{t}^{t_{0}}\}_{t\geq t_{0}}$.
$u_{s,t}\colon\Omega\times\mathbb{R}^{d}\to\mathbb{R}^{d},(\omega,x)\mapsto u_{s,t}(\omega,x):=X_{t}^{s}(\omega,x)$ (2.10)
satisfies the semiflow property
$u_{s,t}(x)=u_{r,t}(u_{s,r}(x)),\quad s\leq r\leq t,~{}x\in\mathbb{R}^{d},~{}\text{almost surely}.$ (2.11)
To establish the existence and uniqueness of random periodic solution of (1.1), several auxiliary lemmas are stated and proved below. The first one considers the uniform boundedness for the exact solution of (1.1).
@@
Applying the Itô formula shows that for any $t_{0}\in\mathbb{R}$, $t\geq t_{0}$ and $p\in(2,p^{*}]$,
$e^{p\lambda t}|X_{t}^{t_{0}}|^{p}=$ $~{}e^{p\lambda t_{0}}|\xi|^{p}+\int_{t_{0}}^{t}p\lambda e^{p\lambda s}|X_{s}^{t_{0}}|^{p}+pe^{p\lambda s}|X_{s}^{t_{0}}|^{p-2}\big{\langle}X_{s}^{t_{0}},-AX_{s}^{t_{0}}\big{\rangle}$
$~{}+pe^{p\lambda s}|X_{s}^{t_{0}}|^{p-2}\big{\langle}X_{s}^{t_{0}},f(s,X_{s}^{t_{0}})\big{\rangle}+\frac{p}{2}e^{p\lambda s}|X_{s}^{t_{0}}|^{p-2}|g(s,X_{s}^{t_{0}})|^{2}$
$~{}+\frac{p(p-2)}{2}e^{p\lambda s}|X_{s}^{t_{0}}|^{p-4}|(X_{s}^{t_{0}})^{\top}g(s,X_{s}^{t_{0}})|^{2}{\,\rm{d}}{s}$
$~{}+\int_{t_{0}}^{t}pe^{p\lambda s}|X_{s}^{t_{0}}|^{p-2}\big{\langle}X_{s}^{t_{0}},g(s,X_{s}^{t_{0}}){\,\rm{d}}{W_{s}}\big{\rangle}.$
@@
$~{}\mathbb{E}\big{[}e^{p\lambda t}|X_{t}^{t_{0}}|^{p}\big{]}$
$\leq$ $~{}\mathbb{E}\big{[}e^{p\lambda t_{0}}|\xi|^{p}\big{]}+\int_{t_{0}}^{t}pe^{p\lambda s}\mathbb{E}\big{[}|X_{s}^{t_{0}}|^{p-2}\big{(}\lambda|X_{s}^{t_{0}}|^{2}+\big{\langle}X_{s}^{t_{0}},-AX_{s}^{t_{0}}\big{\rangle}\big{)}\big{]}{\,\rm{d}}{s}$
$~{}+\int_{t_{0}}^{t}pe^{p\lambda s}\mathbb{E}\Big{[}|X_{s}^{t_{0}}|^{p-2}\Big{(}\big{\langle}X_{s}^{t_{0}},f(s,X_{s}^{t_{0}})\big{\rangle}+\frac{p-1}{2}|g(s,X_{s}^{t_{0}})|^{2}\Big{)}\Big{]}{\,\rm{d}}{s}$
$\leq$ $~{}\mathbb{E}\big{[}e^{p\lambda t_{0}}|\xi|^{p}\big{]}+\int_{t_{0}}^{t}pe^{p\lambda s}\big{(}\widetilde{L_{f}}\mathbb{E}\big{[}|X_{s}^{t_{0}}|^{p}\big{]}+\widetilde{C}\mathbb{E}\big{[}|X_{s}^{t_{0}}|^{p-2}\big{]}\big{)}{\,\rm{d}}{s}.$
@@
$\widetilde{C}|X_{s}^{t_{0}}|^{p-2}=$ $~{}\frac{p}{p-2}\times\bigg{(}(\lambda-\widetilde{L_{f}}-\varepsilon)^{\frac{p-2}{p}}|X_{s}^{t_{0}}|^{p-2}\times\frac{p-2}{p}(\lambda-\widetilde{L_{f}}-\varepsilon)^{\frac{2-p}{p}}\widetilde{C}\bigg{)}$
$\leq$ $~{}\frac{p}{p-2}\times\bigg{(}\frac{p-2}{p}(\lambda-\widetilde{L_{f}}-\varepsilon)|X_{s}^{t_{0}}|^{p}+\frac{2}{p}\Big{(}\frac{p-2}{p}(\lambda-\widetilde{L_{f}}-\varepsilon)^{\frac{2-p}{p}}\widetilde{C}\Big{)}^{\frac{p}{2}}\bigg{)}$
$\leq$ $~{}(\lambda-\widetilde{L_{f}}-\varepsilon)|X_{s}^{t_{0}}|^{p}+\frac{2}{p-2}\big{(}(\lambda-\widetilde{L_{f}}-\varepsilon)^{\frac{2-p}{p}}\widetilde{C}\big{)}^{\frac{p}{2}},$
@@
$\mathbb{E}\big{[}e^{p\lambda t}|X_{t}^{t_{0}}|^{p}\big{]}\leq\mathbb{E}\big{[}e^{p\lambda t_{0}}|\xi|^{p}\big{]}+Ce^{p\lambda t}+\int_{t_{0}}^{t}p(\lambda-\varepsilon)\mathbb{E}\big{[}e^{p\lambda s}|X_{s}^{t_{0}}|^{p}\big{]}{\,\rm{d}}{s}.$
@@
$\mathbb{E}\big{[}e^{p\lambda t}|X_{t}^{t_{0}}|^{p}\big{]}\leq$ $~{}\mathbb{E}\big{[}e^{p\lambda t_{0}}|\xi|^{p}\big{]}+Ce^{p\lambda t}+\int_{t_{0}}^{t}p(\lambda-\varepsilon)\big{(}\mathbb{E}\big{[}e^{p\lambda t_{0}}|\xi|^{p}\big{]}+Ce^{p\lambda s}\big{)}e^{\int_{s}^{t}p(\lambda-\varepsilon){\,\rm{d}}{u}}{\,\rm{d}}{s}$
$\leq$ $~{}\mathbb{E}\big{[}e^{p\lambda t_{0}}|\xi|^{p}\big{]}+Ce^{p\lambda t}+\mathbb{E}\big{[}e^{p\lambda t_{0}}|\xi|^{p}\big{]}e^{p(\lambda-\varepsilon)(t-t_{0})}+C\frac{\lambda-\varepsilon}{\varepsilon}e^{p\lambda t}.$
@@
Let $X_{t}^{t_{0}}(\xi)$ and $X_{t}^{t_{0}}(\eta)$ be two solutions of (1.1) with different initial values $\xi$ and $\eta$, respectively. Suppose that Assumption 2.1 holds for both initial values $\xi$ and $\eta$. Then for any $p\in[2,p^{*}],t_{0}\in\mathbb{R}$ and $t\geq t_{0}$,
$\mathbb{E}\big{[}|X_{t}^{t_{0}}(\xi)-X_{t}^{t_{0}}(\eta)|^{p}\big{]}\leq e^{p(L_{f}-\lambda)(t-t_{0})}\mathbb{E}\big{[}|\xi-\eta|^{p}\big{]}.$ (2.13)
@@
$~{}e^{p\lambda t}|X_{t}^{t_{0}}(\xi)-X_{t}^{t_{0}}(\eta)|^{p}$
$=$ $~{}e^{p\lambda t_{0}}|\xi-\eta|^{p}+\int_{t_{0}}^{t}p\lambda e^{p\lambda s}|X_{s}^{t_{0}}(\xi)-X_{s}^{t_{0}}(\eta)|^{p}$
$~{}+pe^{p\lambda s}|X_{s}^{t_{0}}(\xi)-X_{s}^{t_{0}}(\eta)|^{p-2}\big{\langle}X_{s}^{t_{0}}(\xi)-X_{s}^{t_{0}}(\eta),-A(X_{s}^{t_{0}}(\xi)-X_{s}^{t_{0}}(\eta))\big{\rangle}{\,\rm{d}}{s}$
$~{}+\int_{t_{0}}^{t}pe^{p\lambda s}|X_{s}^{t_{0}}(\xi)-X_{s}^{t_{0}}(\eta)|^{p-2}\big{\langle}X_{s}^{t_{0}}(\xi)-X_{s}^{t_{0}}(\eta),f(s,X_{s}^{t_{0}}(\xi))-f(s,X_{s}^{t_{0}}(\eta))\big{\rangle}$
$~{}+\frac{p}{2}e^{p\lambda s}|X_{s}^{t_{0}}(\xi)-X_{s}^{t_{0}}(\eta)|^{p-2}|g(s,X_{s}^{t_{0}}(\xi))-g(s,X_{s}^{t_{0}}(\eta))|^{2}$
$~{}+\frac{p(p-2)}{2}e^{p\lambda s}|X_{s}^{t_{0}}(\xi)-X_{s}^{t_{0}}(\eta)|^{p-4}|(X_{s}^{t_{0}}(\xi)-X_{s}^{t_{0}}(\eta))^{\top}(g(s,X_{s}^{t_{0}}(\xi))-g(s,X_{s}^{t_{0}}(\eta)))|^{2}{\,\rm{d}}{s}$
$~{}+\int_{t_{0}}^{t}pe^{p\lambda s}|X_{s}^{t_{0}}(\xi)-X_{s}^{t_{0}}(\eta)|^{p-2}\big{\langle}X_{s}^{t_{0}}(\xi)-X_{s}^{t_{0}}(\eta),(g(s,X_{s}^{t_{0}}(\xi))-g(s,X_{s}^{t_{0}}(\eta))){\,\rm{d}}{W_{s}}\big{\rangle}.$
@@
$\mathbb{E}\big{[}e^{p\lambda t}|X_{t}^{t_{0}}(\xi)-X_{t}^{t_{0}}(\eta)|^{p}\big{]}\leq$ $~{}\mathbb{E}\big{[}e^{p\lambda t_{0}}|\xi-\eta|^{p}\big{]}+\int_{t_{0}}^{t}p\mathbb{E}\Big{[}e^{p\lambda s}|X_{s}^{t_{0}}(\xi)-X_{s}^{t_{0}}(\eta)|^{p-2}$
×$~{}\times\Big{(}\big{\langle}X_{s}^{t_{0}}(\xi)-X_{s}^{t_{0}}(\eta),f(s,X_{s}^{t_{0}}(\xi))-f(s,X_{s}^{t_{0}}(\eta))\big{\rangle}$
$~{}+\frac{p-1}{2}|g(s,X_{s}^{t_{0}}(\xi))-g(s,X_{s}^{t_{0}}(\eta))|^{2}\Big{)}\Big{]}{\,\rm{d}}{s}$
$\leq$ $~{}\mathbb{E}\big{[}e^{p\lambda t_{0}}|\xi-\eta|^{p}\big{]}+\int_{t_{0}}^{t}pL_{f}\mathbb{E}\big{[}e^{p\lambda s}|X_{s}^{t_{0}}(\xi)-X_{s}^{t_{0}}(\eta)|^{p}\big{]}{\,\rm{d}}{s}.$
@@
$\mathbb{E}\big{[}e^{p\lambda t}|X_{t}^{t_{0}}(\xi)-X_{t}^{t_{0}}(\eta)|^{p}\big{]}\leq$ $~{}\mathbb{E}\big{[}e^{p\lambda t_{0}}|\xi-\eta|^{p}\big{]}e^{pL_{f}(t-t_{0})},$